OpenStudy (anonymous):

A common resistance of carbon resistors is 33 Ω . How many 33 Ω resistor, connected in parallel, are needed to give a total resistance of 5.5 Ω ?

OpenStudy (mathmate):

Resistors connectors connected in parallel are given by the formula: 1/R=(1/R1+1/R2+1/R3....+1/Rn) if each of n resistors are 33\(\Omega\), then the equation becomes 1/R=(1/33+1/33+....+1/33) .............(n times) =n/33.......................(1) Since the final resistance needs to be 5.5\(\Omega\), substitute R=5.5 in equation (1) and solve for n.